suppose there are 1000 (type a) people who value flood control more than the 1000 (type b) people.type a demand qd = 100 − p type b demand qd = 50 − p where q measures the quality of flood control. if the price of a unit of flood control is $100,000 and the citizens of toadhop did not work together, what would be the amount of flood control purchased? (2 marks) 


1
Expert's answer
2021-06-16T04:26:17-0400

If the price of a unit of flood control is $100,000 and the citizens of toadhop did not work together, then the amount of flood control purchased would be:

"qd = 1,000\u00d7(100 - p) + 1,000\u00d7(50 - p) = 150,000 - 2,000\u00d7100 = -50,000."